OpenStudy (sambot):

A cube with an edge of length s has a volume of 64 units. What is the length of the edge?

jhonyy9 (jhonyy9):

formula for volume of a cube do you know what is ?

OpenStudy (sambot):

No..

jhonyy9 (jhonyy9):

hi Sam ! i dont see you here much time

jhonyy9 (jhonyy9):

hope you remember me

OpenStudy (sambot):

Yeah I do

jhonyy9 (jhonyy9):

ok. ty so the volume of a cube you get multiplied area of base time the height so what will be ? do you get it from this ?

jhonyy9 (jhonyy9):

the base is a square area of a square what will be ?

OpenStudy (sambot):

I think it is length times width

jhonyy9 (jhonyy9):

yes liike a case of a rectangle but hope you know that a square has every sides length equal - yes ?

OpenStudy (sambot):

Yes

jhonyy9 (jhonyy9):

so in this case what will be with side length s ?

jhonyy9 (jhonyy9):

area square = s*s = ?

OpenStudy (sambot):

Umm...I'm confused

jhonyy9 (jhonyy9):

= s^2 yes ?

jhonyy9 (jhonyy9):

so and you know that the length of the height of a cube is the length of a side what is s too yes ? so than we know that the volume of a cube is equal area of the base time the height so what will be ?

jhonyy9 (jhonyy9):

V = area of the base *height V = s^2 * height height = the length of an edge = s V = s^2 * ?

OpenStudy (sambot):

Ok..

jhonyy9 (jhonyy9):

do you get the length of an edge than you know that the volume is equal 64 units cubed ?

jhonyy9 (jhonyy9):

V = s^2 *s = ? 64 = s^?

jhonyy9 (jhonyy9):

what will be the exponent of s than you multiply s squared by s ?

jhonyy9 (jhonyy9):

V = s^2 *s = s^(2+1) = ?

OpenStudy (sambot):

I'm really confused

jhonyy9 (jhonyy9):

V = s^3 so the volume of a cube equal the length of edge cubed what mean that you need getting a number what cubed will give you 64 on other words s = cuberoot(64) = ?

jhonyy9 (jhonyy9):

can you getting this number ?

OpenStudy (sambot):

So do I have to find the square root of 64?

jhonyy9 (jhonyy9):

not sqrt64 bc. there above you get s cubed equal the volume

OpenStudy (sambot):

Oh

OpenStudy (sambot):

Oh

jhonyy9 (jhonyy9):

not square root you need getting cube root of 64 http://www.calculatorsoup.com/calculators/algebra/cuberoots.php try there please

jhonyy9 (jhonyy9):

do you can opening this site please ?

OpenStudy (sambot):

It says the answer is 4

jhonyy9 (jhonyy9):

there is a cube root calculator where you just write there this 64 and push the button of calcule yes this is equal exactly 4

jhonyy9 (jhonyy9):

bc. 4*4*4 = 16*4 = 64

jhonyy9 (jhonyy9):

so do you understand it now ?

OpenStudy (sambot):

Yes I do

jhonyy9 (jhonyy9):

and how many will be than the length of an edge of this cube ?

jhonyy9 (jhonyy9):

what is the number what you ve get with this calculator ?

OpenStudy (sambot):

Um..the only number I got with this question was 64

jhonyy9 (jhonyy9):

yes and using this site with this calculator of cube root so you get for cube root 64 = ?

jhonyy9 (jhonyy9):

so was 4 what mean that the length of an edge equal 4 units ok. ?

OpenStudy (firechild17):

@3mar @Will.H @brooke..help00

OpenStudy (sambot):

4?

jhonyy9 (jhonyy9):

hope helped yes sure 4 units

jhonyy9 (jhonyy9):

bc. cube root 64 equal 4

OpenStudy (sambot):

Ok.. thanks

OpenStudy (firechild17):

@MissMeow @Slenderman

jhonyy9 (jhonyy9):

good luck and anytime bye bye

OpenStudy (firechild17):

@Afro

OpenStudy (firechild17):

@jabez177

OpenStudy (firechild17):

@Hayhayz

OpenStudy (firechild17):

@Nnesha

OpenStudy (firechild17):

Guys help her plz

OpenStudy (will.h):

The question is solved and closed. Why r u tagging?

OpenStudy (afro):

dude that guy answered it so nice

OpenStudy (afro):

ikr he put it piece by piece in the best way possible

jabez177 (jabez177):

Yep! It's easy.

OpenStudy (3mar):

Sorry! I was no here.